# How can Diffusion Models Evolve into Continual Generators?

### v2

@@ 3.2 Basic Diffusion Model Training
3.3 Inter-task Knowledge Consistency
In CDG, the inter-task knowledge consistency term, formulated as $\mathcal{L}_{IKC}=\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}$, serves as a critical metric to quantify the divergence between score approximators across tasks. Unlike continual classification, where L2-regularization techniques [29] may offer moderate performance improvements, their direct application in generative settings can induce catastrophic degradation, severely impairing the model’s generative capability, as evidenced in Tab. 1. To circumvent this limitation, we introduce a new knowledge retention strategy, drawing inspiration from [8], where the previously learned score estimator $\epsilon_{\theta}^{p}$ acts as a teacher to guide the adaptation of $\epsilon_{\theta}^{q}$. In contrast to conventional knowledge distillation techniques [8], which primarily manipulate class probability distributions, our approach capitalizes on the stochastic gradients governed by the reverse-time SDE. This sophisticated formulation not only facilitates an exceptionally seamless and cohesive knowledge retention but also profoundly mitigates the propensity for CF, thereby preserving generative fidelity across successive tasks.
Let $\mathcal{M}_{k-1}$ denote the frozen diffusion model for task $\mathcal{T}_{k-1}$, parameterized by $\theta_{k-1}$. For a new task $\mathcal{T}_{k}$, we seek to adapt $\theta_{k}$ while preserving the score-matching capability on prior tasks. To achieve this, we minimize the Bregman divergence [27] between the score distributions of $\mathcal{M}_{k-1}$ and $\mathcal{M}_{k}$ over a shared noise manifold. Specifically, given the current samples $(x_{t}^{k},y^{k})\sim p^{k}$ from task $\mathcal{T}_{k}$ and the replayed samples $(\hat{x}_{t}^{k},\hat{y}^{k})\sim\mathcal{B}_{k}$, the $\mathcal{L}_{IKC}$ is defined as:
@@ 3.3 Inter-task Knowledge Consistency
$\begin{split}\varphi(\hat{x}_{t}^{k},\hat{y}^{k},t)=\mathbb{E}\Bigl{[}\nabla_{\hat{x}_{t}^{k}}\log\epsilon_{\theta_{k-1}}(\hat{x}_{t}^{k}|\hat{y}^{k},t)\times\nabla_{\hat{x}_{t}^{k}}\log\epsilon_{\theta_{k-1}}(\hat{x}_{t}^{k}|\hat{y}^{k},t)^{\top}\Bigr{]}.\end{split}$ (9)
@@
3.4 Unconditional Knowledge Consistency
Building upon the inter-task model alignment established in Section 3.3, we now operationalize the $\mathcal{L}_{UKC}$ term from Theorem 4, which governs consistency in the mean of unconditional sample embeddings and reverse-time denoising trajectories. This component bridges theoretical bounds with practical implementation by explicitly aligning the mean functions of reverse processes across tasks.
In CDG, deriving task-specific reverse mean functions presents a fundamental challenge. Given current instances $(x_{0}^{k},y^{k})\sim(p^{k}\cup\mathcal{B}_{k})$  [^5] In the UKC term, the primary objective is to preserve geometric constraints and retain knowledge across both historical and current data. In the ensuing subsections, we reinforce this continuity by integrating historical data, thereby treating the current samples as a comprehensive amalgamation of both past and present information. and buffered historical samples $(\hat{x}_{0}^{k},\hat{y}^{k})\sim\mathcal{B}_{k}$, direct computation of $\mathbb{E}[x_{t}]$ remains ill-posed due to the artificial noise-driven construction of $x_{t}$ in diffusion frameworks. This arises from the intrinsic semantic mismatch in perturbed diffused states $x_{t}$ across training phases, rendering naive trajectory averaging and constraint imposition ineffective for gradient-based optimization. To address this, we devise an indirect mean constraint through symbiotic knowledge transfer between the frozen teacher $\mathcal{M}_{k-1}$ and adaptive student $\mathcal{M}_{k}$. By reparameterizing $\mathbb{E}[x_{t-1}]$, we enable full gradient backpropagation while enforcing coherent diffused space constraints. Crucially, label-marginalized computation ensures these constraints govern unconditional generation fidelity. The reverse process mean functions for both historical and current models are derived analytically within the shared data using their respective noise prediction networks:
@@ 3.4 Unconditional Knowledge Consistency
Here, $\mu_{\theta_{k-1}}(\cdot)$ and $\mu_{\theta_{k}}(\cdot)$ are the analytically back-calculated means prescribed by historical and current models.
@@
The weighting term $\frac{\bar{\alpha}_{t}^{2}}{1-\bar{\alpha}_{t}^{2}}$ emphasizes alignment during semantically critical mid-diffusion phases, where latent structures transition between noise and meaningful representations. By penalizing deviations in denoising trajectories and the mean of unconditional sample embeddings, this loss enforces constraints that preserve the manifold topology of historical data within the evolving student model $\mathcal{M}_{k}$. This mechanism complements the instantaneous model matching strategy outlined in Section 3.3, ensuring both local knowledge coherence and global structural fidelity.
3.5 Label Knowledge Consistency
The label knowledge consistency term, denoted as $\mathcal{L}_{LKC}$, aims to ensure consistency in the shared semantic part of the label space across original samples. To effectively operationalize this concept, we introduce a label regressor, which is not used for optimizing label classification performance, but rather to explore the semantic similarity between $\hat{x}_{0}$ and $x_{0}$ in the label space, similar to multi-domain alignment methods [10, 21, 15, 5, 47] in zero-shot learning. Moreover, this approach effectively avoids the singularity of label classification, and can be similarly applied when the prior knowledge in diffusion models is in the form of text or other modalities. Let $\mathcal{M}_{k}$ represent the current model for task $\mathcal{T}_{k}$, with the corresponding label regressor $h_{\phi}^{k}$ that adapts to the task-specific distribution. The label regressor from the previous task, $\mathcal{M}_{k-1}$, is frozen to retain the semantic structure of earlier tasks and is represented by the label regressor $h_{\phi}^{<k}$. Within this framework, we ideally generate samples from $p_{t}(x_{0}|x_{t})$, i.e., the previous task, to obtain $\hat{x}^{\prime}_{0}=\mathcal{M}_{k-1}(\epsilon,\hat{y},T)$. However, such sampling incurs substantial computational overhead and resource consumption. Moreover, in $\mathcal{M}_{k-1}$, the generated images are often nearly indistinguishable from the original ones, rendering explicit sampling redundant. Thus, for efficiency considerations, we directly sample $\hat{x}_{0}$ from $\mathcal{B}_{k}$ to serve as the generated samples, which are then used alongside the current samples $x_{0}$ to compute $\mathcal{L}_{LKC}$.
@@ 3.5 Label Knowledge Consistency
$\mathcal{L}_{LKC}=\frac{\bar{\alpha}_{t}}{\bar{\beta}_{t}}\mathbb{E}_{\hat{x}_{0}\sim\mathcal{B}_{k}}\left[D_{\mathrm{KL}}\left(h_{\phi}^{<k}(y|\hat{x}_{0})\parallel h_{\phi}^{k}(y|x_{0})\right)\right].$ (12)
@@
This objective promotes semantic consistency between historical and current samples by extracting shared prior knowledge, thereby mitigating drift within the label space. Ultimately, through the integration of all three optimization strategies, diffusion models achieve continuous alignment in model parameters, unconditional mean distributions, and the label space, effectively preserving shared knowledge across continual tasks, thereby achieving improved long-term generative performance.
4 Experiments
4.1 Datasets and Benchmark Characteristics
@@
We conduct comprehensive experiments on four representative vision benchmarks spanning diverse domains. MNIST [14] consists of 60,000 training and 10,000 test images spanning 10 handwritten digit classes. In our setting, the dataset is partitioned into 5 disjoint tasks, each comprising a subset of the digit classes. CIFAR-100 [13] comprises 100 object categories, each containing 600 images. The dataset poses significant challenges for CDG due to its low image resolution and minimal knowledge overlap across tasks. It is partitioned into 10 tasks, each with 10 classes. Flowers102 [23] consists of 8,189 images from 102 flower species with significant intra-class variance. It is divided into 10 tasks for fine-grained generation evaluation. CUB-200-2011 [36] offers 11,788 bird images from 200 species with subtle morphological variations, split into 10 tasks. To ensure fairness and maintain consistent GPU memory consumption, all images are uniformly resized to $32\times 32$.
4.2 Metrics
To establish a rigorous theoretical framework for evaluating CF in CDG, we formulate two orthogonally motivated metrics: Mean Fidelity (MF) as a terminal-state performance quantifier and Incremental Mean Fidelity (IMF) as a long-term evaluation quantifier. Given a temporally ordered task manifold $\{\mathcal{T}_{k}\}_{k=1}^{K}$, MF interrogates cross-task generative performance through $\text{MF}=\mathbb{E}_{k\sim[1,K]}[d_{\mathcal{M}}(p^{k}_{\text{real}}\parallel p^{k}_{\text{gen}})]$, where $d_{\mathcal{M}}$ operationalizes a generative quality metric (e.g., FID [7]) between ground-truth and synthesized distributions. Conversely, IMF induces a time-incremental aggregation mechanism: $\text{IMF}=\mathbb{E}_{k\sim[1,K]}[\mathbb{E}_{i\leq k}[d_{\mathcal{M}}(p^{i}_{\text{real}}\parallel p^{i}_{\text{gen}})]]$, formally encoding the expected trajectory stability under sequential updates. This dual-metric paradigm dissects both terminal performance degradation (via MF’s terminal-state evaluation) and streaming generation robustness (through IMF’s incremental analysis). To eliminate confounding factors, we enforce strict experimental controls including fixed-output cardinality (2048 samples) and balanced class-conditioned sampling, ensuring metrological consistency. The proposed metrics establish a comprehensive evaluation protocol for assessing both instance-specific generation fidelity and longitudinal knowledge preservation in non-stationary generative scenarios.
4.3 Experimental Implementation
All investigations are rigorously conducted under identical hardware setups and environmental conditions to eliminate confounding variables. The neural architecture leverages the diffusers library-implemented UNet2DModel with hierarchical residual connections, featuring four stages of downsampling/upsampling blocks (128→256→256→256 channels) where the third downblock and second upblock integrate attention mechanisms. The label regressor is strategically deployed in the penultimate feature space to mitigate interference from raw noise patterns. A uniform learning rate of 1e-3 is systematically applied across all trials, coupled with the Adam optimizer for optimization. To ensure an absolutely equitable experimental framework, identical training epochs, batch sizes, and dataset partitioning strategies are meticulously maintained for each benchmark dataset. The complete experimental protocol is executed on four NVIDIA A100-40G GPU accelerators. For a more detailed model architecture diagram, please refer to Appendix C.
Table 1: A comprehensive performance comparison across various buffer configurations and methodologies is conducted, ensuring that all experiments are executed under identical hardware setups and environmental conditions to maintain fairness and reproducibility.
Method MNIST-5T CIFAR100-10T Flowers102-10T CUB200-10T Average
MF ($\downarrow$) IMF ($\downarrow$) MF ($\downarrow$) IMF ($\downarrow$) MF ($\downarrow$) IMF ($\downarrow$) MF ($\downarrow$) IMF ($\downarrow$) MF ($\downarrow$) IMF ($\downarrow$)
Non-CL 4.99 4.99 65.97 65.97 21.16 21.16 35.80 35.80 31.98 31.98
Rehearsal-free Methods
KD 62.83 58.98 114.22 103.40 102.49 117.79 157.70 111.07 109.31 97.81
EWC 119.29 176.59 110.07 141.78 104.91 114.35 165.19 163.08 124.86 148.95
L2 105.34 67.66 134.25 129.48 153.53 132.70 125.47 108.61 129.65 109.61
SI 85.39 66.95 145.06 132.67 148.00 136.80 142.90 132.01 130.34 117.11
MAS 291.70 304.40 146.77 169.19 104.99 317.51 128.66 154.84 168.03 236.48
C-LoRA 106.71 80.63 134.58 131.47 369.62 363.88 142.85 129.43 188.44 176.35
Storage Rehearsal Methods (buffer=512)
ER 97.57 69.38 94.19 89.30 88.36 94.06 130.57 135.07 102.67 96.95
A-GEM 94.67 66.00 96.85 90.80 89.41 92.59 121.73 139.66 100.66 97.26
CCD (Ours) 79.76 57.63 96.34 90.14 86.11 92.18 121.15 101.42 95.84 85.34
Storage Rehearsal Methods (buffer=2560)
ER 73.60 56.95 103.38 98.55 67.89 80.33 373.50 353.32 154.59 147.29
A-GEM 69.77 54.32 104.58 99.43 66.36 79.79 103.78 114.93 86.12 87.12
CCD (Ours) 67.48 50.94 101.51 98.68 65.31 79.70 106.93 84.11 85.31 78.36
Storage Rehearsal Methods (buffer=5120)
ER 60.40 135.19 103.34 99.39 67.19 80.23 135.98 164.87 91.73 119.92
A-GEM 68.59 107.24 103.87 100.49 65.39 79.71 70.33 99.21 77.05 96.66
CCD (Ours) 48.97 40.08 102.09 98.33 64.97 79.62 67.00 61.37 70.76 69.85
4.4 Performance Comparison
In Tab. 1, we integrate classification-inspired CL methods into the CDG setting, executing a comprehensive study of their continual generative efficacy. Our evaluation spans established CL methods including Knowledge Distillation (KD) [8], Elastic Weight Consolidation (EWC) [12], L2 Regularization [29], Synaptic Intelligence (SI) [45], Memory-Aware Synapses (MAS) [2], Continual-LoRA (C-LoRA) [30], Experience Replay (ER) [18], and Averaged-Gradient Episodic Memory (A-GEM) [4]. Interestingly, our evaluation reveals that even stable diffusion models suffer from generative collapse under standard CL regimes, resembling GAN instabilities, particularly on the CUB and Flowers102 benchmarks. This underscores the urgent need for a foundational re-evaluation of CDG methods before advancing novel techniques. To this end, we systematically dissect classical paradigms, establishing essential baselines for this emerging research frontier.
Experimental results demonstrate that our method achieves substantial performance gains on the MNIST and CUB200 datasets, with the IMF metric showing marked superiority over ER [18] and A-GEM [4] across all buffer size configurations. This strong empirical evidence confirms that our CCD framework enables stable and high-quality continual generation in CDG, particularly on datasets with high inter-task knowledge overlap, while effectively mitigating the CF phenomenon. On the Flowers102 dataset, which contains relatively few training samples and permits the buffer to store a representative portion of the data, our method maintains a consistent, though moderate, advantage over baseline approaches. However, the performance gap is less pronounced compared to MNIST and CUB200. As anticipated, results on the CIFAR100 dataset are less favorable, with some metrics underperforming relative to traditional ER [18]. This limitation likely stems from the low inter-task knowledge correlation inherent in coarse-grained datasets, where minimal shared knowledge across tasks renders aggressive retention strategies counterproductive, potentially impairing future generative performance. This finding highlights a key open challenge for future research: developing effective strategies for knowledge retention and propagation under conditions of minimal cross-task overlap.
Collectively, our CCD framework achieves state-of-the-art performance across three benchmark datasets, excluding CIFAR100, delivering substantial average improvements in both MF and IMF metrics under all three buffer size configurations. Moreover, the training process remains notably stable, with no observed instances of cross-task generative collapse. These results offer foundational insights and practical guidance for advancing continual data generation, particularly in addressing the stability challenges unique to sequential generative model training. To further illustrate the effectiveness of our approach, we present visualizations of samples generated for the first task at the final training stage in Appendix D. These qualitative results provide a more intuitive and direct assessment of generative retention compared to numerical metrics.

### v3

@@ Introduction
Inter-task Knowledge Consistency
The inter-task knowledge consistency loss $\mathcal{L}_{IKC}=\epsilon_{\theta}^{q}-\epsilon_{\theta}^{p}$ measures, for each sample $x$ and diffusion step $t$, the output and parameter difference between the estimator trained on task $q$ and the estimator retained from task $p$, thereby directly capturing their divergence across tasks. In CC tasks, L2 regularization is commonly employed to prevent excessive variations in parameters and outputs. However, in CDG, this approach may lead to catastrophic degradation, severely impairing the model’s generative capability, as demonstrated in Table CDG Pipeline. To circumvent this limitation, we introduce a new knowledge retention strategy, drawing inspiration from (Hinton 2015), where the previously learned score estimator $\epsilon_{\theta}^{p}$ acts as a teacher to guide the adaptation of $\epsilon_{\theta}^{q}$. In contrast to conventional knowledge distillation techniques (Moslemi et al. 2024), which primarily manipulate class probability distributions, our approach capitalizes on the stochastic gradients governed by the reverse-time SDE. This sophisticated formulation not only facilitates an exceptionally seamless and cohesive knowledge retention but also profoundly mitigates the propensity for GCF, thereby preserving generative fidelity across successive tasks.
Let $\mathcal{M}_{k-1}$ denote the frozen diffusion model for task $\mathcal{T}_{k-1}$, parameterized by $\theta_{k-1}$. For a new task $\mathcal{T}_{k}$, we seek to adapt $\theta_{k}$ while preserving the score-matching capability on prior tasks. To achieve this, we minimize the Bregman divergence (Siahkamari et al. 2020) between the score distributions of $\mathcal{M}_{k-1}$ and $\mathcal{M}_{k}$ over a shared noise manifold. Specifically, given the current samples $(x_{t}^{k},y^{k})\sim p^{k}$ from task $\mathcal{T}_{k}$ and the replayed samples $(\hat{x}_{t}^{k},\hat{y}^{k})\sim\mathcal{B}_{k}$, the $\mathcal{L}_{IKC}$ is defined as:
@@
$\begin{split}\varphi(\hat{x}_{t}^{k},\hat{y}^{k},t)=\mathbb{E}\Bigl{[}\nabla_{\hat{x}_{t}^{k}}\log\epsilon_{\theta_{k-1}}(\hat{x}_{t}^{k}|\hat{y}^{k},t)\,\times\\ \nabla_{\hat{x}_{t}^{k}}\log\epsilon_{\theta_{k-1}}(\hat{x}_{t}^{k}|\hat{y}^{k},t)^{\top}\Bigr{]}.\end{split}$ (9)
@@
Unconditional Knowledge Consistency
Building on the inter-task model alignment, we now instantiate the $\mathcal{L}_{UKC}$ term from Theorem 1, which enforces consistency in the mean of unconditional sample embeddings and reverse-time denoising trajectories. This component serves as a bridge between theoretical guarantees and practical implementation by explicitly aligning the mean functions of reverse processes across tasks.
Deriving task-specific reverse mean functions presents a fundamental challenge. Given current instances $(x_{0}^{k},y^{k})\sim(p^{k}\cup\mathcal{B}_{k})$  [^5] In $\mathcal{L}_{UKC}$, the primary objective is to preserve geometric constraints and retain knowledge across both historical and current data. In the ensuing subsections, we reinforce this continuity by integrating historical data, thereby treating the current samples as a comprehensive amalgamation of both past and present information. and buffered historical samples $(\hat{x}_{0}^{k},\hat{y}^{k})\sim\mathcal{B}_{k}$, direct computation of $\mathbb{E}[x_{t}]$ remains ill-posed due to the artificial noise-driven construction of $x_{t}$ in diffusion frameworks. This arises from the intrinsic semantic mismatch in perturbed diffused states $x_{t}$ across training phases, rendering naive trajectory averaging and constraint imposition ineffective for gradient-based optimization. To address this, we devise an indirect mean constraint through symbiotic knowledge transfer between the frozen teacher $\mathcal{M}_{k-1}$ and adaptive student $\mathcal{M}_{k}$. By reparameterizing $\mathbb{E}[x_{t-1}]$, we enable full gradient backpropagation while enforcing coherent diffused space constraints. Crucially, label-marginalized computation ensures these constraints govern unconditional generation fidelity. The reverse process mean functions for both historical and current models are derived analytically within the shared data using their respective noise prediction networks:
@@
Here, $\mu_{\theta_{k-1}}(\cdot)$ and $\mu_{\theta_{k}}(\cdot)$ are the posterior means predicted by the previous and current models, respectively. $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ are standard DDPM forward coefficients.
@@
The weighting term $\frac{\bar{\alpha}_{t}^{2}}{1-\bar{\alpha}_{t}^{2}}$ emphasizes alignment during semantically critical mid-diffusion phases, where latent structures transition between noise and meaningful representations. By penalizing deviations in denoising trajectories and the mean of unconditional sample embeddings, this loss enforces constraints that preserve the manifold topology of historical data within the evolving student model $\mathcal{M}_{k}$. This mechanism complements the instantaneous model matching strategy outlined in $\mathcal{L}_{\mathrm{IKC}}$, ensuring both local knowledge coherence and global structural fidelity.
Prior Knowledge Consistency
The prior knowledge consistency loss $\mathcal{L}_{PKC}$ preserves the shared semantic prior, instantiated as label information in this work, across the original samples. We realize it with a label regressor that gauges the semantic proximity of $\hat{x}_{0}$ and $x_{0}$ in label space, following multi-domain alignment strategies in zero-shot learning (Hwang and Sigal 2014; Ni, Zhang, and Xie 2019; Li et al. 2023; Duan et al. 2024; Zhang et al. 2024b). Because the regressor is trained only for similarity, not classification accuracy, it avoids label-collapse and readily generalizes to textual or other modality priors in diffusion models. Let $\mathcal{M}_{k}$ denote the current task model with its task-adaptive regressor $h_{\phi}^{k}$, and let the frozen regressor $h_{\phi}^{<k}$ from $\mathcal{M}_{k-1}$ preserve earlier semantics. Although one could generate $\hat{x}^{\prime}_{0}=\mathcal{M}_{k-1}(\epsilon,\hat{y},T)$ via $p_{t}(x_{0}|x_{t})$, this generative replay is costly and produces images almost identical to stored ones. Hence, we simply draw $\hat{x}_{0}$ from the buffer $\mathcal{B}_{k}$ and pair it with the current sample $x_{0}$ to compute $\mathcal{L}_{PKC}$.
@@
$\mathcal{L}_{PKC}=\frac{\bar{\alpha}_{t}}{\bar{\beta}_{t}}\mathbb{E}_{\hat{x}_{0}\sim\mathcal{B}_{k}}\left[D_{\mathrm{KL}}\left(h_{\phi}^{<k}(y|\hat{x}_{0})\parallel h_{\phi}^{k}(y|x_{0})\right)\right].$ (12)
@@
This objective enforces semantic consistency between past and current samples by extracting shared prior knowledge, thereby curbing label-space drift. Ultimately, together with the other two optimizations, diffusion models achieve continuous alignment in model parameters, unconditional mean distributions, and the prior space, effectively preserving shared knowledge across continual generation tasks, thereby achieving improved long-term generative performance.
Experiments
Datasets and Benchmark Characteristics
@@
We conduct comprehensive experiments on five representative vision benchmarks spanning diverse domains. MNIST (LeCun et al. 1998) consists of 60,000 training and 10,000 test images spanning 10 handwritten digit classes. In our setting, the dataset is partitioned into 5 disjoint tasks, each comprising a subset of the digit classes. OxfordPets (Parkhi et al. 2012) comprises 7,349 RGB images across 37 classes. To accommodate the CDG pipeline, we split it into five tasks, excluding two classes to maintain an equal number of categories per task. CIFAR-100 (Krizhevsky and Hinton 2009) comprises 100 object categories, each containing 600 images. The dataset poses significant challenges for our CCD optimization framework due to its low image resolution and minimal knowledge overlap across tasks. It is partitioned into 10 tasks, each with 10 classes. Flowers102 (Nilsback and Zisserman 2008) consists of 8,189 images from 102 flower species with significant intra-class variance. It is divided into 10 tasks for fine-grained generation evaluation. CUB-200-2011 (Wah et al. 2011) offers 11,788 bird images from 200 species with subtle morphological variations, split into 10 tasks. To ensure fairness and consistent GPU memory consumption, all images are uniformly resized to $32\times 32$.
CDG Pipeline
We present an end-to-end pipeline as a rigorous and standardized framework for diffusion-based continual generation methods. Centered on a unified UNet2D diffusion backbone (see Appendix Figure. 6), the model is trained using the standard DDPM (Ho, Jain, and Abbeel 2020) formulation with 1000 denoising steps and an MSE loss derived from Gaussian noise prediction. Optimization employs the Adam optimizer with a batch size of 200 and a fixed learning rate of $1\times 10^{-3}$. During inference, a configurable DDIM (Song, Meng, and Ermon 2020) scheduler with 50 sampling steps is used. For evaluation, 2048 samples are generated and uniformly distributed across classes to ensure metric fairness. All experiments are conducted on NVIDIA A800 GPUs under consistent computational settings. To rigorously assess GCF, we introduce two complementary metrics: Mean Fidelity (MF), capturing terminal-state generative quality, and Incremental Mean Fidelity (IMF), quantifying temporal stability. Given a temporally ordered task manifold $\{\mathcal{T}_{k}\}_{k=1}^{K}$, MF is defined as $\mathrm{MF}=\mathbb{E}_{k\sim[1,K]}[d_{\mathcal{M}}(p^{k}_{\text{real}}\parallel p^{k}_{\text{gen}})]$, where $d_{\mathcal{M}}$ denotes a generative quality metric (e.g., FID (Heusel et al. 2017)). IMF extends this by aggregating performance over time: $\mathrm{IMF}=\mathbb{E}_{k\sim[1,K]}[\mathbb{E}_{i\leq k}[d_{\mathcal{M}}(p^{i}_{\text{real}}\parallel p^{i}_{\text{gen}})]]$, reflecting expected stability under continual updates. This dual-metric formulation enables precise characterization of both endpoint fidelity and generative consistency across task sequences.
[…]
MF↓ IMF↓ MF↓ IMF↓ MF↓ IMF↓ MF↓ IMF↓ MF↓ IMF↓ MF↓ IMF↓
Non-CL 4.99 4.99 224.26 224.26 65.97 65.97 21.16 21.16 35.80 35.80 1.00 1.00
[…]
LwF TPAMI (2017) 62.83 58.98 288.35 283.91 114.22 103.40 102.49 117.79 157.70 111.07 4.97 4.66
EWC ICLR (2017) 119.29 176.59 307.85 333.72 110.07 141.78 104.91 114.35 165.19 163.08 7.30 9.80
L2 105.34 67.66 272.00 267.30 134.25 129.48 153.53 132.70 125.47 108.61 7.02 5.20
SI PMLR (2017) 85.39 66.95 269.22 267.75 145.06 132.67 148.00 136.80 142.90 132.01 6.30 5.35
MAS ECCV (2018) 291.70 304.40 201.77 222.17 146.77 169.19 104.99 317.51 128.66 154.84 14.03 16.78
C-LoRA TMLR (2024) 106.71 80.63 457.49 442.48 134.58 131.47 369.62 363.88 142.85 129.43 9.38 8.19
[…]
ER NeurIPS (2017) 97.57 69.38 287.12 276.05 94.19 89.30 88.36 94.06 130.57 135.07 6.02 4.94
A-GEM ICLR (2019) 94.67 66.00 289.42 279.36 96.85 90.80 89.41 92.59 121.73 139.66 5.87 4.83
LwF + ER - 89.46 105.41 306.58 290.70 112.80 113.97 245.86 319.55 119.10 152.15 7.19 8.70
DCM CVPR (2024) 63.06 43.74 290.77 282.36 100.06 96.91 187.25 259.39 112.62 143.58 5.49 5.55
[…]
ER NeurIPS (2017) 73.60 56.95 288.19 280.73 103.38 98.55 67.89 80.33 373.50 353.32 6.25 5.56
A-GEM ICLR (2019) 69.77 54.32 287.59 280.86 104.58 99.43 66.36 79.79 103.78 114.93 4.58 4.13
LwF + ER - 75.42 115.63 296.24 286.72 106.63 107.52 212.12 316.77 95.64 144.92 6.15 9.02
DCM CVPR (2024) 61.11 108.93 285.79 281.81 96.24 89.15 199.57 264.15 107.00 138.27 5.48 8.16
[…]
ER NeurIPS (2017) 60.40 135.19 276.00 274.99 103.34 99.39 67.19 80.23 135.98 164.87 4.38 7.64
A-GEM ICLR (2019) 68.59 107.24 272.63 273.84 103.87 100.49 65.39 79.71 70.33 99.21 4.32 6.04
LwF + ER - 57.25 72.87 277.31 278.84 106.13 97.61 212.12 316.77 88.88 142.35 5.37 7.25
DCM CVPR (2024) 50.45 25.48 285.79 279.52 101.62 100.79 199.57 264.15 110.10 119.37 5.09 4.74
[…]
Table 1: Performance comparison across datasets and buffer sizes. Top two per buffer size are marked in bold and underline. Weighted Averages are computed as the normalized mean of MF↓ and IMF↓ using the Non-CL scores as the baseline.
Performance Comparison
In this subsection, we benchmark representative baselines under standardized hardware and software settings. These include CC methods, LwF (Li and Hoiem 2017), EWC (Kirkpatrick et al. 2017), SI (Zenke, Poole, and Ganguli 2017), MAS (Aljundi et al. 2018), ER (Lopez-Paz and Ranzato 2017), and A-GEM (Chaudhry et al. 2019), as well as generative CL approaches such as C-LoRA (Smith et al. 2024) and DCM (Ye and Bors 2024). The implementation details are provided in the Appendix. Although buffer-based CL methods like DDGR (Gao and Liu 2023), TD (Li et al. 2025), AM (Wan et al. 2025), and CUTER (Wang et al. 2025) perform well in CC tasks, they rely on sample embeddings or prototypes, an assumption incompatible with diffusion models, whose latent space consists of isotropic Gaussian noise. Therefore, we do not consider them in this work. We also exclude approaches such as (Zając et al. 2023; Masip et al. 2023; Cheng et al. 2024) that depend on pretrained diffusion backbones and generative replay buffers, as they violate the controlled comparison protocol. These exclusions highlight the methodological gap between continual classification and diffusion-based generation, motivating the treatment of our CDG pipeline as a standalone research subject.
To ensure fair comparison, we apply our CCD optimization framework atop the ER buffer. As shown in Table CDG Pipeline, CCD consistently outperforms most baselines across varying buffer sizes and evaluation metrics, with its advantage becoming more pronounced as memory increases. This trend is especially evident on MNIST-5T, Flowers102-10T, and CUB200-10T, where CCD reduces MF and IMF scores by over 50% relative to ER in some cases, indicating substantial mitigation of GCF. Compared to the state-of-the-art DCM, CCD demonstrates superior resilience to catastrophic collapse, particularly on fine-grained datasets such as Flowers102-10T and CUB200-10T, where it achieves lower IMF scores. Moreover, on OxfordPets-5T, where most buffer-based methods fail to yield any meaningful generative improvement, our method, at a buffer size of 2560, nearly matches the performance of a Non–CL upper bound. Finally, while DCM performance saturates with larger buffers, CCD continues to drive metrics lower, highlighting its scalability.
Nevertheless, we identify a key limitation: since CCD is designed to preserve shared knowledge across tasks, its effectiveness diminishes on coarse-grained datasets like CIFAR100-10T, where semantic overlap is minimal. This finding highlights a key open challenge for future research: developing effective strategies for knowledge retention and propagation under conditions of minimal cross-task overlap.